OpenStudy (anonymous):

If f(x) = ax^2 + bx + a is divided by (x – b) a , b ≠ 0 , remainder is ‘a’ , then a is (A) b^2 (B) -1 (C) 0 (D) a + b Explain

hartnn (hartnn):

remainder when f(x) = ax^2 + bx + a is divided by (x – b), is just f(b) so find f(b) first

hartnn (hartnn):

then equate it to 'a' f(b)=a

OpenStudy (anonymous):

by using remainder theorem f(b)=a*b^2+b*b+a=(a+1)b^2+a=a[given] a+1=0------>a=-1:)